Today is Lara's birthday. She turned 24. Lara is twice as old as Joe was when Lara was as old as Joe is now. How old is Joe now?
Answer by Edwin McCravy(8880)

Today is Lara's birthday. She turned 24. Lara is twice as old as Joe was when Lara was as old as Joe is now. How old is Joe now?
Suppose Lara is x years older than Joe. 

So x years ago, Lara was Joe's age now.  

She was 24-x then, and that's what Joe's age is now.

So Joe's age now = 24-x 

So Joe's age then = (24-x)-x ---- or 24-x-x or 24-2x

Lara's age now = 2 times Joe's age then

            24 = 2(24-2x)
            24 = 48 - 4x
            4x = 24
             x = 6

So Joe's age now is 24-x or 24-6 or 18.

Checking:  Lara is 24.  6 years ago, Lara was 18 and Joe was 12.  
And sure enough 24 is twice 12. So her age now is twice his age then.

Edwin